OpenStudy (anonymous):

solve. -2/7+6+1/7x=18

OpenStudy (campbell_st):

I'd start by simplifying \[5\frac{5}{7} + \frac{1}{7} x = 18\] subtract 5 5/7 from both sides gives \[\frac{1}{7}x = 12\frac{2}{7}\] then multiply every term by 7 to find x

OpenStudy (anonymous):

i still dont get it @campbell_st